OpenStudy (anonymous):

perpendicular to y=-1/5x+6 passing through the point (5,-6) THE EQUATION OF THE LINE IN SLOPE INTERCEPT FORM IS?

OpenStudy (anonymous):

\[slope_{The unknown line }=\frac{ -1 }{ \frac{ -1 }{ 5 } }=5\] \[The ..equation..of ..the..line..is..\frac{ y+6 }{ x-5 }=5....y+6=5(x-5)\]

OpenStudy (anonymous):

slope of the given line = -1/5 slope of the line perpendicular to the given line can be found from the formula  find the slope and substitute in the eqn y = mx+c substitute x and y as (5,-6) to get c value